OpenStudy (anonymous):

@MathematicYou have to use ONe to One property (1/5)^X+1=125 @MathematicYou have to use ONe to One property (1/5)^X+1=125 @Mathematics

OpenStudy (anonymous):

is that \[(\frac{1}{5})^{x+1} \]?

OpenStudy (anonymous):

or \[(\frac{1}{5})^x +1 \]

jhonyy9 (jhonyy9):

\[5^{-(x+1)}=5^{3}\]

jhonyy9 (jhonyy9):

so than -x-1=3 x=-4